The amount of time students study for a final exam is normally distributed with a mean of 25 hours and a standard deviation of 4.0 hours. What is the probability that a student will study for more than 26 hours?

a).2500

b).4013

c).5987

d) .6554
Answer #1

Solution :

Given that,

mean = = 25

standard deviation = = 4

P(x >26 ) = 1 - P(x< 26)

= 1 - P[(x -) / < (26-25) /4 ]

= 1 - P(z <0.25 )

Using z table

= 1 -  0.5987

probability= 0.4013